OpenStudy (anonymous):

Find the volume of the region under the graph of f(x,y) = 5x+y+1 and above the region y^2\le x, 0\le x\le 4.

OpenStudy (anonymous):

@myininaya the second part is \[y^2 \le x, 0 \le x \le 4\]

OpenStudy (matlee):

myninia i took ap statistics and they kicked me out cause statitics test is in may

OpenStudy (anonymous):

so i think id set it up as\[\int\limits_{0}^{4} \int\limits_{?}^{?} 5x+y+1 dy dx\]

OpenStudy (anonymous):

i just dont know what the y bounds would be if that is correct. one might be y=x^2

OpenStudy (anonymous):

The region \(y^2\leq x, 0\leq x\leq 4\)?

OpenStudy (anonymous):

That is a weird region

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

I think I would use \(y\leq \sqrt x\).

OpenStudy (anonymous):

The top limit will give \(y=-5x-1\).

OpenStudy (anonymous):

It doesn't make sense, because the top part is below the bottom part.

OpenStudy (anonymous):

Hmm, wait up, this is volume so it is a triple integral...

OpenStudy (anonymous):

double integral

OpenStudy (anonymous):

Maybe use:\[\int\limits_{0}^{4} \int\limits_{0}^{\sqrt x} 5x+y+1~ dy dx\] I don't like how they say "above the region". I think they mean within the region. I also don't like how there is no seeming lower limit for \(y\).

OpenStudy (anonymous):

naa that didnt work

OpenStudy (anonymous):

On other thing you could try then:\[\int\limits_{0}^{4} \int\limits_{-\sqrt x}^{\sqrt x} 5x+y+1~ dy dx\]

OpenStudy (anonymous):

swag

myininaya (myininaya):

did that work @pmkat14 ? just curious myself

OpenStudy (anonymous):

yea!!!!